OpenStudy (anonymous):

Write the standard equation for the circle center (–6, 7), r = 9.

jimthompson5910 (jim_thompson5910):

(x-h)^2+(y-k)^2 = r^2 (x-(-6))^2+(y-7)^2 = 9^2 (x+6)^2+(y-7)^2 = 81

OpenStudy (anonymous):

can you help me with 2 more?

jimthompson5910 (jim_thompson5910):

sure I can help you out

OpenStudy (anonymous):

Write the standard equation for the circle center (–6, –8), that passes through (0, 0).

jimthompson5910 (jim_thompson5910):

distance from (-6,-8) to (0,0) is the radius d = sqrt((0-(-6))^2+(0-(-8))^2) d = sqrt((0+6)^2+(0+8)^2) d = sqrt((6)^2+(8)^2) d = sqrt(36+64) d = sqrt(100) d = 10 So exact distance between the two points is 10 units. This means that the radius is 10. So r = 10 The center is (-6,-8), so h = -6 and k = -8 (x-h)^2+(y-k)^2 = r^2 (x-(-6))^2+(y-(-8))^2 = 10^2 (x+6)^2+(y+8)^2 = 100 So the equation of the circle is (x+6)^2+(y+8)^2 = 100

OpenStudy (anonymous):

The center of a circle is (h, 7) and the radius is 10. The circle passes through (3, –1). Find all possible values of h. i think the answer is 9,-3

jimthompson5910 (jim_thompson5910):

You are 100% correct. Nice job.

OpenStudy (anonymous):

ahhhhhh thank you sooooooooo much!!!!!!

jimthompson5910 (jim_thompson5910):

you're welcome